OpenStudy (anonymous):

How to solve 1 = (1/2 + 2/3)t I used, w = rt

mathslover (mathslover):

We have to solve for t?

OpenStudy (anonymous):

mhmm

mathslover (mathslover):

\[\large{1=(\frac{1}{2}+\frac{2}{3})\times t}\] \[\large{1=(\frac{3+4}{6})\times t}\] \[\large{1=\frac{7}{6}\times t}\] \[\large{\frac{6}{7}=t}\] got it?

OpenStudy (anonymous):

Ohh I see thanks! :D